OpenStudy (anonymous):

The height h(t), in feet, of an airborne tee-shirt t seconds after being launched can be approximated by h(t)= -15t^2+120t+10, 0

OpenStudy (anonymous):

so you need to find when H(t) equals 115 and at what time does it happen?

OpenStudy (anonymous):

\[115=-15t ^{2}+120t+10\] solve thsi

OpenStudy (anonymous):

ok yes i do

OpenStudy (anonymous):

ok i will try now

OpenStudy (anonymous):

I have -15t^2+120t+105=0

OpenStudy (anonymous):

i subtract 115-10=105

OpenStudy (anonymous):

should be -15t^2+120t-105=0 because you do 10-115=-105

OpenStudy (anonymous):

oh ok I missed the negative sign for 105

OpenStudy (anonymous):

ok this is how far i have gotten and i am not sure what to do next could you guide me?

OpenStudy (anonymous):

do i need to know what two numbers multiply for 120t?

OpenStudy (anonymous):

might be best to use the quadratic formula

OpenStudy (anonymous):

well this how it suppose to look i think -15t^2(t-?)(t-?)

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] where a=-15 b=120 and c=-105

OpenStudy (anonymous):

ok so how you set up the math problem on here where do you go to find those type of setup?

OpenStudy (anonymous):

equation button

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i see where 4ac you said c=105 is placed but what do i do with 4ac

OpenStudy (anonymous):

it should look something like this\[\frac{ -120\pm \sqrt{120^{2}-4\times-15\times-105} }{ 2\times-15 }\]

OpenStudy (anonymous):

oh ok i will try to solve this difficult problem